OpenStudy (anonymous):

5. Which ordered pair is a solution of the equation y = –5x + 4? (1 point) (2, 14) (5, –29) (3, –19) (5, –21)

OpenStudy (anonymous):

just plug in the pairs and check the answer, for example : if we use the last answer (5,-21) first number belongs to x second to y so (-21)=-5(5)+4 -21=-25+4 -21=-21 it is true so that is your answer.

OpenStudy (anonymous):

(-)(-)=+ (-)(+)=- (+)(+)=+ (+)(-)=- Just in case.